OpenStudy (anonymous):

can you show me how to get the least coomon multilple of the denominator for 2/3+1/6

OpenStudy (anonymous):

When you add fractions with the same denominator, it's very easy. 1/4 + 2/4 = 3/4. 3/9 +4/9 = 7/9. 8/10213 + 11/10213 = 19/10213. You'd add the numerators normally, and the denominators would be the same. And it makes sense that a quarter slice of pizza plus 2 quarter slices give you 3 quarter slices of the pizza. But sometimes, you need to add fractions that don't have common denominators. Like 1/4 + 1/2. Or 1/3 + 4/9. What you can do then is to "common denomintorize" them. What is the least common multiple of 2 and 4? Multiples of 2 are 2, 4, 6, 8, 10, ... Multiples of 4 are 4, 8, 12, 16, ... What is the smallest number that's in both lists? Well that'd be 4. Or the LCM(2, 4) = 4. That means we want to make the fractions in 1/4 + 1/2 have a denominator of 4. So we'd like to express 1/2 as a fraction with a denominator of 4. We can multiply 1/2 on the numerator and denominator by 2 to get a denominator 4. giving us that 1/2 = 2/4. So now 1/4 + 1/2 = 1/4 + 2/4. And that's just 3/4. So in your problem what's the least common multiple of 3 and 6?